OpenStudy (cantfindasymptotewithbothhands):

12x+13x=14 I'm asked to solve by completing the square (but we can solve the best/easiest way). First I multiply both sides by 1/12 so the coefficient of x = 1. I am getting confused when I divide the second term by two and square. 13/12x becomes 13/24^2 becomes 169/576. Just doesn't look right to me... any thoughts?

OpenStudy (anonymous):

Well, I think you're over-complicating it! :D I would combine the like terms, 12x & 13x. So then, you have 25x=14. Divide each side by 25, and you get 0.56 or 14/25.

OpenStudy (cantfindasymptotewithbothhands):

oh, my bad. that first term is 12x^2 :)

OpenStudy (anonymous):

Oh, okay! So, you want the right side to equal to 0, so subtract 14 from both sides. That would give you \[12^2 + 13x - 14 = 0\] Factor it. \[(-7 -4x)(2 -3x) = 0\] So, in order for the equation to come out to be 0, one of the binomials needs to equal 0, right? So either (-7-4x) equals 0, or (2-3x) equals 0. Let's solve the first one. -7-4x = 0 Add 7 to both sides. -4x = 7 Then, divide each side by -4. 7/ -4 is -1.75 Now we have our first answer. Let's go back to the second binomial, and solve it. 2-3x = 0 Subtract 2 from both sides. -3x = -2 Divide each side by -3. -2/-3 is about 0.67. So x is equal to either -1.75 OR 0.67. Does that make sense?

OpenStudy (cantfindasymptotewithbothhands):

I don't see from which stage you started factoring, from the original equation: 12x^2+13x-14=0 or from the point after I multiplied by 12/1 to clear the coefficient of x^2: x^2+13/12x-14/12=0

OpenStudy (cantfindasymptotewithbothhands):

(by the way, my name is Dana as well) :D

OpenStudy (anonymous):

I started factoring from the original equation\[12x^2+13x-14=0\] You have a fantastic name :D

OpenStudy (cantfindasymptotewithbothhands):

I get it now :) For some reason my brain starts to have buffer overflow errors when I try to factor coefficients of x^ when it's not prime :) Thank you!

OpenStudy (anonymous):

You're welcome! Glad I could help :D